OpenStudy (anonymous):

What is the x-intercept of the graph y = -4x + 8? -4 0 2 8

OpenStudy (anonymous):

8

OpenStudy (anonymous):

put x=0 y=8 8 is x intercept

OpenStudy (diyadiya):

x intercept is when y=0 0=-4x + 8 4x =8 x=8/4=2

OpenStudy (anonymous):

no no just look to be y=ax+or-b

OpenStudy (anonymous):

look to b

OpenStudy (diyadiya):

@moha_10 that is y-intercept